OpenStudy (anonymous):

The root of a radical becomes the denominator when the expression is rewritten as a rational exponent. always sometimes never

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

if we have\[\sqrt{25}\] written as \[25^{1/2}\] I do not believe this is in the denominator?

OpenStudy (turingtest):

Actually it is true. They are saying\[\sqrt[3]{x}\to x^{1/3}\]notice that the 3 in the root went into the denominator of the exponent. That is the denominator they are referring to. The rule in general is\[\sqrt[b]{x^a}=x^{a/b}\]her the root was called b and as you can see, it too went into the denominator.